# Efficient Online Learning and Adaptive Planning for Robotic Information Gathering Based on Streaming Data

### v1

@@ Abstract
Keywords - Informative path planning, Gaussian processes, robotic information gathering, active mapping.
@@ III-C Adaptive planning
6:         Perform SSGP regression with new batch $X,Y$
@@
14:Find $\mathcal{S}=\arg\max_{\nu}I(\nu)$ s.t $\nu\in\mathbb{T}$ and compute_distance$(\nu_{0},\mathcal{S})<=r$
@@
16:$\mathcal{P}\leftarrow\texttt{extract\_path}(\nu_{0},\mathcal{S})$
17:$\nu_{0}\leftarrow\mathcal{S}$
@@ IV-A Simulation results
It can be observed from Table I that our method is as accurate as the RIG-tree method in mapping the environment in terms of the RMSE. It is also observed that our method traverses longer distances than the RIG-tree, due to our replanning strategy that replans paths on the go. SSGP regression is also as accurate as mapping as compared to exact GP regression. Fig. 3 shows a plot of the progression of the root mean squared error (RMSE) and the mean entropy, calculated on a discretized grid of size $100$x100100100100, for the various planning strategies with SSGP regression. Our planning strategy achieves the lowest entropy in the predictions compared to the other methods, which could be attributed to the robot traversing larger distances with our method.
@@
Our method ($\epsilon=0.1$) $0.18\pm 0.04$1.21 $\pm$ 0.23 $599.31\pm 12.08$
Our method ($\epsilon=0.2$) $0.19\pm 0.04$ $-1.12\pm 0.41$ 600.57 $\pm$ 13.17
Our method ($\epsilon=0.3$) $0.17\pm 0.02$ $-0.93\pm 0.39$ $597.37\pm 9.08$
RIG-tree Re.h (GP) $0.20\pm 0.04$ $-1.11\pm 0.16$ $551.82\pm 3.34$
RIG-tree Re.h (SGP) $0.24\pm 0.02$ $-0.06\pm 0.06$ $553.45\pm 28.4$
RIG-tree Re.h (SSGP) 0.14 $\pm$ 0.02 $-1.16\pm 0.30$ $557.89\pm 1.85$
Random planner (GP) $0.28\pm 0.08$ $-0.79\pm 0.21$ $560.21\pm 9.54$
@@
Random planner (SSGP) $0.28\pm 0.05$ $-0.54\pm 0.25$ $555.67\pm 5.82$

### v2

@@ III-C Adaptive planning
6:   Optimize hyperparameters with SSGP regression for new batch $X,Y$
@@
[…]
@@
[…]
@@ IV-A Simulation results
It can be observed from Table I that our method is as accurate as the RIG-tree method in mapping the environment in terms of the RMSE. It is also observed that our method traverses longer distances than the RIG-tree, due to our replanning strategy that replans paths on the go. SSGP regression is also as accurate as mapping as compared to exact GP regression. Fig. 3 shows a plot of the progression of the root mean squared error (RMSE) and the mean entropy, calculated on a discretized grid of size 100100x100100, for the various planning strategies with SSGP regression. Our planning strategy achieves the lowest entropy in the predictions compared to the other methods, which could be attributed to the robot traversing larger distances with our method.
@@
Our method ($\epsilon=0.1$) $0.18\pm 0.04$ -1.21 $\pm$ 0.23 $599.31\pm 12.08$
Our method ($\epsilon=0.2$) $0.19\pm 0.04$ $1.12\pm 0.41$ 600.57 $\pm$ 13.17
Our method ($\epsilon=0.3$) $0.17\pm 0.02$ $0.93\pm 0.39$ $597.37\pm 9.08$
RIG-tree Re.h (GP) $0.20\pm 0.04$ $1.11\pm 0.16$ $551.82\pm 3.34$
RIG-tree Re.h (SGP) $24\pm 0.02 0.0$$06\pm 0.06$ $553.45\pm 28.4$
RIG-tree Re.h (SSGP) 0.14 $\pm$ 0.02 $1.16\pm 0.30$ $557.89\pm 1.85$
Random planner (GP) $0.28\pm 0.08$ $0.79\pm 0.21$ $560.21\pm 9.54$
@@
Random planner (SSGP) $0.28\pm 0.05$ $0.54\pm 0.25$ $555.67\pm 5.82$